At 1:00 p.m., a car leaves a city and travels north at a rate of 55 mi/h. An hour later a second car leaves the city and travels south at a rate of 60 mi/hr. At what time will the two cars be 285 miles apart?
Answer by checkley75(3666)

55t+60(t-1)=285
55t+60t-60=285
115t=285+60
115t=345
t=345/115
t=3 hours after the first car leaves they will be 285 miles apart
proof
55*3+60*2=285
165+120=285
285=285